OpenStudy (anonymous):

How does the order of operations relate to solving multi-step equations? Use examples of solving an equation, and then evaluating it to help in your explanation.

OpenStudy (anonymous):

Multi-step equations must be solved using the order of operations or they will give multiple different answers. Here's an example. 2(x+1)=6 Go out of order and subtract the 1 first. Perhaps it would look something like this. 2(x+1)=6 -1 -1 2x = 5 Then, divide by 2 on each side. I am going to multiply by 1/2 since I don't have a good way to show a division line. (1/2)(2x)=5(1/2) x=5/2 or x=2.5 Correctly, the solution does the multiplication required by the parentheses on the left first. 2(x+1)=6 2x+2=6 Now, we can subtract the two from each side. 2x+2=6 -2 -2 2x=4 Now divide by 2 on each side. I will multiply by 1/2 for ease of typing. (1/2)(2x)=4(1/2) x=2 Both 2 and 2.5 cannot be the solution. The order-of-operation rules allow every person to get the same answer and let math be "true."